# Mora: Enabling Generalist Video Generation via A Multi-Agent Framework

### v2

@@ 3.1 Agent-based Video Generation
Definition and Specialization of Agents.   The definition of agents enables flexibility in the breakdown of complex work into smaller and more specific tasks. Solving different video generation tasks often requires the collaboration of agents with diverse abilities, each contributing specialized outputs. In our framework, we have 5 basic roles: prompt selection and generation agent, text-to-image generation agent, image-to-image generation agent, and image-to-video generation agent and video-to-video agent.
Figure 2: Illustration of how to use Mora to conduct video-related tasks.
  • Prompt Selection and Generation Agent: Prior to the commencement of the initial image generation, textual prompts undergo a rigorous processing and optimization phase. This critical agent can employ large language models like GPT-4, Llama Achiam et al. (2023); Touvron et al. (2023). It is designed to meticulously analyze the text, extracting pivotal information and actions delineated within, thereby significantly enhancing the relevance and quality of the resultant images. This step ensures that the textual descriptions are thoroughly prepared for an efficient and effective translation into visual representations.
  • Text-to-Image Generation Agent: The text-to-image model Rombach et al. (2022b); Podell et al. (2023) stands at the forefront of translating these enriched textual descriptions into high-quality initial images. Its core functionality revolves around a deep understanding and visualization of complex textual inputs, enabling it to craft detailed and accurate visual counterparts to the provided textual descriptions.
  • Image-to-Image Generation Agent: This agent Brooks et al. (2023) works to modify a given source image in response to specific textual instructions. The core of its functionality lies in its ability to interpret detailed textual prompts with high accuracy and subsequently apply these insights to manipulate the source image accordingly. This involves a detailed recognition of the text’s intent, translating these instructions into visual modifications that can range from subtle alterations to transformative changes. The agent leverages a pre-trained model to bridge the gap between textual description and visual representation, enabling seamless integration of new elements, adjustment of visual styles, or alteration of compositional aspects within the image.
  • Image-to-Video Generation Agent  Blattmann et al. (2023): Following the creation of the initial image, the Video Generation Model is responsible for transitioning the static frame into a vibrant video sequence. This component delves into the analysis of both the content and style of the initial image, serving as the foundation for generating subsequent frames. These frames are meticulously crafted to ensure a seamless narrative flow, resulting in a coherent video that upholds temporal stability and visual consistency throughout. This process highlights the model’s capability to not only understand and replicate the initial image but also to anticipate and execute logical progressions in the scene.
  • Video Connection Agent: Utilizing the Video-to-Video Agent, we create seamless transition videos based on two input videos provided by users. This advanced agent selectively leverages key frames from each input video to ensure a smooth and visually consistent transition between them. It is designed with the capability to accurately identify the common elements and styles across the two videos, thus ensuring a coherent and visually appealing output. This method not only improves the seamless flow between different video segments but also retains the distinct styles of each segment.
Every agent is responsible for the specific input and output. These results can be utilized for different designed tasks.
Approaches.   By setting the agents’ roles and operational skills, we can define basic workflows for different tasks. We design six text-to-video generation tasks: (1) Text-to-video generation, (2) Text-conditional image-to-video generation, (3) Extend generated videos, (4) Video-to-video editing, (5) Connect Videos, and (6) Simulate digital worlds, which are described below.
  • Task 1: Text-to-Video Generation: This task harnesses a detailed textual prompt from the user as the foundation for video creation. The prompt must meticulously detail the envisioned scene. Utilizing this prompt, the Text-to-Image agent utilizes the text, distilling themes and visual details to craft an initial frame. Building upon this foundation, the Image-to-Video component methodically generates a sequence of images. This sequence dynamically evolves to embody the prompt’s described actions or scenes, and each video is derived from the last frame from the previous video, thereby achieving a seamless transition throughout the video.
  • Task 2: Text-conditional Image-to-Video Generation: Task 2 mirrors the operational pipeline of Task 1, with a key distinction. Unlike Task 1 with only texts as inputs, Task 2 integrates both a textual prompt and an initial image into the Text-to-Image agent’s input. This dual-input approach enriches the content generation process, enabling a more nuanced interpretation of the user’s vision.
  • Task 3: Extend Generated Videos: This task focuses on extending the narrative of an existing video sequence. By taking the last frame of an input video as the starting point, the video generation agent crafts a series of new, coherent frames that continue the story. This approach allows for the seamless expansion of video content, creating longer narratives that maintain the consistency and flow of the original sequence.
  • Task 4: Video-to-Video Editing: Task 4 introduces a sophisticated editing capability, leveraging both the Image-to-Image and Image-to-Video agents. The process begins with the Image-to-Image agent, which takes the first frame of an input video and applies edits based on the user’s prompt, achieving the desired modifications. This edited frame then serves as the initial image for the Image-to-Video agent, which generates a new video sequence that reflects the requested obvious or subtle changes, offering a powerful tool for dynamic video editing.
  • Task 5: Connect Videos: The Image-to-Video agent leverages the final frame of the first input video and the initial frame of the second input video to create a seamless transition, producing a new video that smoothly connects the two original videos.
  • Task 6: Simulating Digital Worlds: This task specializes in the whole style changing for video sequences set in digitally styled worlds. By appending the phrase "In digital world style" to the edit prompt, the user instructs the Image-to-Video agent to craft a sequence that embodies the aesthetics and dynamics of a digital realm or utilize the Image-to-Image agent to transfer the real image to digital style. This task pushes the boundaries of video generation, enabling the creation of immersive digital environments that offer a unique visual experience.
3.2 Implementation Detail of Agents
Prompt Selection and Generation.  Currently, GPT-4  Achiam et al. (2023) stands as the most advanced generative model available. By harnessing the capabilities of GPT-4, we are able to generate and meticulously select high-quality prompts. These prompts are detailed and rich in information, facilitating the Text-to-Image generation process by providing the agent with comprehensive guidance.
Text-to-Image Generation.  We utilize the pretrained large text-to-image model to generate a high-quality and representative first image.
The Stable Diffusion XL (SDXL) Podell et al. (2023) is utilized for the first implementation. It introduces a significant evolution in the architecture and methodology of latent diffusion models Rombach et al. (2022b); Meng et al. (2022) for text-to-image synthesis, setting a new benchmark in the field. At the core of its architecture is an enlarged UNet backbone Ronneberger et al. (2015) that is three times larger than those used in previous versions of Stable Diffusion 2 Rombach et al. (2022b). This expansion is principally achieved through an increased number of attention blocks and a broader cross-attention context, facilitated by integrating a dual text encoder system. The first encoder is based on OpenCLIP Ilharco et al. (2021) ViT-bigG Cherti et al. (2023); Radford et al. (2021); Schuhmann et al. (2022), while the second utilizes CLIP ViT-L, allowing for a richer, more nuanced interpretation of textual inputs by concatenating the outputs of these encoders. This architectural innovation is complemented by the introduction of several novel conditioning schemes that do not require external supervision, enhancing the model’s flexibility and capability to generate images across multiple aspect ratios. Moreover, SDXL features a refinement model that employs a post-hoc image-to-image transformation to elevate the visual quality of the generated images. This refinement process utilizes a noising-denoising technique, further polishing the output images without compromising the efficiency or speed of the generation process.
Image-to-Image Generation.  Our initial framework realize uitilze InstructPix2Pix as Image-to-Image generation agent. InstructPix2Pix  Brooks et al. (2023) are intricately designed to enable effective image editing from natural language instructions. At its core, the system integrates the expansive knowledge of two pre-trained models: GPT-3  Brown et al. (2020) for generating editing instructions and edited captions from textual descriptions, and Stable Diffusion  Rombach et al. (2022a) for transforming these text-based inputs into visual outputs. This ingenious approach begins with fine-tuning GPT-3 on a curated dataset of image captions and corresponding edit instructions, resulting in a model that can creatively suggest plausible edits and generate modified captions. Following this, the Stable Diffusion model, augmented with the Prompt-to-Prompt technique, generates pairs of images (before and after the edit) based on the captions produced by GPT-3. The conditional diffusion model at the heart of InstructPix2Pix is then trained on this generated dataset. InstructPix2Pix directly utilizes the text instructions and input image to perform the edit in a single forward pass. This efficiency is further enhanced by employing classifier-free guidance for both the image and instruction conditionings, allowing the model to balance fidelity to the original image with adherence to the editing instructions.
Our first implementation utilizes the state-of-the-art video generation model Stable Video Diffusion to generate video. The Stable Video Diffusion (SVD)  Blattmann et al. (2023) architecture introduces a cutting-edge approach to generating high-resolution videos by leveraging the strengths of LDMs Stable Diffusion v2.1  Rombach et al. (2022a), originally developed for image synthesis, and extending their capabilities to handle the temporal complexities inherent in video content. At its core, the SVD model follows a three-stage training regime that begins with text-to-image pertaining, where the model learns robust visual representations from a diverse set of images. This foundation allows the model to understand and generate complex visual patterns and textures. In the second stage, video pretraining, the model is exposed to large amounts of video data, enabling it to learn temporal dynamics and motion patterns by incorporating temporal convolution and attention layers alongside its spatial counterparts. This training is conducted on a systematically curated dataset, ensuring the model learns from high-quality and relevant video content. The final stage, high-quality video finetuning, focuses on refining the model’s ability to generate videos with increased resolution and fidelity, using a smaller but higher-quality dataset. This hierarchical training strategy, complemented by a novel data curation process, allows SVD to excel in producing state-of-the-art text-to-video and image-to-video synthesis with remarkable detail, realism, and coherence over time.
Connect Videos.  For the video connection task, we utilize SEINE Chen et al. (2023a) to connect videos. SEINE is constructed upon a pre-trained diffusion-based T2V model, LaVie Wang et al. (2023a) agent. SEINE centered around a random-mask video diffusion model that generates transitions based on textual descriptions. By integrating images of different scenes with text-based control, SEINE produces transition videos that maintain coherence and visual quality. Additionally, the model can be extended for tasks such as image-to-video animation and autoregressive video prediction.
@@ 4.1 Setup
Baseline.  In the text-to-video generation, existing open-sourced works that show competitive performance are employed as the baseline models, including Videocrafter1 Chen et al. (2023c), Show-1 Zhang et al. (2023a), Pika pik , Gen-2 Gen , ModelScope Mod , LaVie-Interpolation, LaVie Wang et al. (2023b) and CogVideo Hong et al. (2022b). In the other five tasks, we compare Mora with Sora.
Basic Metrics.  For text-to-video generation, we employed several metrics from Vbench Huang et al. (2023) for evaluation from two aspects: video quality and video condition consistency.
For video quality measurement, we use six metrics. ❶ Object Consistency, computed by the DINO Caron et al. (2021) feature similarity across frames to assess whether object appearance remains consistent throughout the whole video; ❷ Background Consistency, calculated by CLIP Radford et al. (2021) feature similarity across frames; ❸ Motion Smoothness, which utilizes the motion priors in the video frame interpolation model AMT Li et al. (2023) to evaluate the smoothness of generated motions; ❹ Aesthetic Score, obtained by using the LAION aesthetic predictor lio on each video frame to evaluate the artistic and beauty value perceived by humans, ❺ Dynamic Degree, computed by employing RAFT Teed and Deng (2020) to estimate the degree of dynamics in synthesized videos; ❻ Imaging Quality, calculated by using MUSIQ Ke et al. (2021) image quality predictor trained on SPAQ Fang et al. (2020) dataset.
For measuring video condition consistency, we use two metrics. ❶ Temporal Style, which is determined by utilizing ViCLIP Wang et al. (2023c) to compute the similarity between video features and temporal style description features, thereby reflecting the consistency of the temporal style; ❷ Appearance Style, by calculating the feature similarity between synthesized frames and the input prompt using CLIP Radford et al. (2021), to gauge the consistency of appearance style.
Self-defined Metrics.  For evaluating other tasks, we also define four metrics.
❶ Video-Text Integration $VideoTI$, devised to enhance the quantitative evaluation of the model’s fidelity to textual instructions. It employs LLaVA Liu et al. (2024b) to transfer input image into textual descriptors $T_{i}$ and Video-Llama Zhang et al. (2023b) to transfer videos generated by the model into textual $T_{v}$. The textual representation of the image is prepended with the original instructional text, forming an augmented textual input $T_{mix}$. Both the newly formed text and the video-generated text will be input to BERT Devlin et al. (2018). The embeddings obtained are analyzed for semantic similarity through the computation of cosine similarity, providing a quantitative measurement of the model’s adherence to the given instructions and image.
$VideoTI=cosine(embed_{mix},embed_{v})$ (1)
❷ Temporal Consistency $TCON$, designed to measure the coherence between an original video and its extended version, provides a vital tool for assessing the integrity of extended video content. For each input-output video pair, we employ ViCLIP Wang et al. (2023c) video encoder to extract their feature vectors. We then compute cosine similarity to get the score.
$TCON=cosine(V_{input},V_{output})$ (2)
❸ Temporal coherence $Tmean$, by quantifying the correlation between the intermediate generated video and the input videos based on $TCON$.
$Tmean=(TCON_{front}+TCON_{beh})/2$ (3)
where $TCON_{front}$ measures the correlation between the intermediate video and the preceding video in the time series while $TCON_{beh}$ assesses the correlation with the subsequent video. The average of these scores provides an aggregate measure of temporal coherence across the video sequence.
❹ Video Length, to evaluate the models’ efficiency in producing video content, specifically focusing on the maximum duration, measured in seconds.

### v3

@@ 3.1 Agent Architecture of Mora
The specialization of agents enables flexibility in the breakdown of complex work into smaller and more focused tasks, as depicted in Figure 1. In our framework, we have five agents: prompt selection and generation agent $A_{1}$, text-to-image generation agent $A_{2}$, image-to-image generation agent $A_{3}$, image-to-video generation agent $A_{4}$, and video-to-video agent $A_{5}$. We present brief descriptions below, and detailed definitions can be found in Appendix A.5.
  • Prompt Selection and Generation Agent ($A_{1}$): This agent employs large language models like GPT-4 and Llama (Achiam et al., 2023; Touvron et al., 2023) to analyze and enhance textual prompts, extracting key information and actions to optimize image relevance and quality.
  • Text-to-Image Generation Agent ($A_{2}$): Utilizing models such as those by Rombach et al. (2022b) and Podell et al. (2023), this agent translates enriched textual descriptions into high-quality images by deeply understanding and visualizing complex inputs.
  • Image-to-Image Generation Agent ($A_{3}$): Referencing Brooks et al. (2023), this agent modifies source images based on detailed textual instructions, accurately interpreting and applying these to make visual alterations ranging from subtle to transformative.
  • Image-to-Video Generation Agent ($A_{4}$): As described by Blattmann et al. (2023), this model transitions static images into dynamic video sequences by analyzing content and style, ensuring temporal stability and visual consistency.
  • Video Connection Agent ($A_{5}$): This agent creates seamless transition videos from two input videos by leveraging keyframes and identifying common elements and styles, ensuring coherent and visually appealing outputs.
General Structure.
Mora model structure is depicted in Figure 1. Specifically, given a user input $T$, the Prompt Enhancement Agent ($A_{1}$) first refines $T$ into a form better suited for video generation. The enhanced prompt is then passed to the Text-to-Image Agent ($A_{2}$) to generate the first frame of the video. At this stage, the user can review and confirm whether the tone and quality of the frame meet their expectations. If not, the user can either request a re-generation or pass the frame to the Image-to-Image Agent ($A_{3}$) for adjustments. Once the desired first frame is finalized, it is forwarded to the Image-to-Video Agent ($A_{4}$) to generate a high-quality video that aligns with the user’s requirements. This step-by-step, independently controllable, and interactive process ensures that the final video more closely meets the user’s expectations while maintaining high quality. In cases where a user wishes to generate a continuous video from different video clips, $A_{5}$ analyzes the final frame of the preceding video and the initial frame of the next and ensures a smooth blending between them. Moreover, this procedural design ensures that the generation process does not have to start from scratch, and the human-in-the-loop technique makes the entire pipeline more controllable, as detailed in Appendix A.4. It enables our framework to handle various tasks, such as image-to-video generation, and even video extension and stitching. For more details about supported tasks, please refer to Appendix A.3. In addition to prompt-based generation, Mora structure also supports task-wise model fine-tuning, ensuring that agents can effectively follow instructions and consistently produce high-quality content.
3.2 Self-Modulated Multi-Agent Finetuning
In this section, we introduce our proposed multi-agent finetuning design, based on the previously described model structure. Directly prompting each agent does not account for the downward transmission of information, which could lead to inefficiencies or errors in communication between agents. Additionally, the impact of each agent on the final result is not uniform. To address these issues, we adopt an end-to-end training approach. Our proposed training procedure involves (1) a self-modulation factor to enhance inter-agent coordination, (2) a data-free training strategy to synthesize training data, (3) LLM selection with human-in-the-loop to control the training data quality. In the following sections, we provide a detailed explanation of each component.
Self-modulated Fine-tuning Algorithm.
Previous methods primarily employ direct end-to-end fine-tuning across the entire task procedure, while others may fine-tune individual agents based on specific tasks. However, both approaches can influence model performance: (1) end-to-end fine-tuning may result in improper loss allocation for each agent, and (2) fine-tuning agents individually can lead to misaligned distributions. To address the limitations of existing fine-tuning approaches, we propose a self-modulated fine-tuning algorithm specifically designed for our multi-agent model structure. This method aims to enhance coordination among agents and improve overall performance by introducing a modulation embedding that dynamically adjusts the influence of each agent during the generation process. The key motivations behind our approach are to balance the impact of each agent on the final output, improve inter-agent communication and coordination, and allow for dynamic adjustment of agent contributions based on the current task and intermediate outputs.
Our design introduces a modulation embedding that is concatenated with the text embedding of the enhanced prompt, allowing for fine-grained control over the generation process. This embedding is optimized alongside the model parameters during training, enabling the system to learn optimal coordination strategies. By doing so, we ensure that each agent can adapt its output based on the state of the preceding agent, leading to more coherent and high-quality video generation.
The implementation of our self-modulated fine-tuning algorithm begins with the initialization of a modulation embedding $\{\mathbf{z}_{i}\}$ using the text embedding of a special token (Ning et al., 2023): $\{\mathbf{z}_{i}=\mathcal{E}_{i}(\text{"{[Mod]}"})\}$] " ) }, where $\mathcal{E}_{i}$ is the text encoder of agent $i$. For each enhanced prompt $P_{\text{enh}}$, we compute its text embedding $\{\mathbf{e}_{i}=\mathcal{E}_{i}(P_{\text{enh}})\}$ and concatenate it with the modulation embedding: $\{\tilde{\mathbf{e}}_{i}=[\mathbf{e}_{i};\mathbf{z}_{i}]\}$. This combined embedding serves as input to the agents.
Each agent $\mathcal{M}_{i}$ in the system processes its input and produces an output $O_{i}$. For the first agent, $O_{1}=\mathcal{M}_{1}(\tilde{\mathbf{e}})$, and for subsequent agents ($i>1$), $O_{i}=\mathcal{M}_{i}(O_{i-1},\tilde{\mathbf{e}})$. The modulation factor, which influences the impact of each agent, is calculated as the L2 norm of $\mathbf{z}_{i}$: $||\mathbf{z}_{i}||_{2}=\sqrt{\sum_{k}z_{k}^{2}}$.
During training, we minimize the total loss $L_{\text{total}}=L(O_{n},V_{\text{target}})$ between the final output $O_{n}$ and the target video $V_{\text{target}}$. Both the model parameters $\theta_{i}$ of each agent and the modulation embedding $\mathbf{z}_{i}$ are updated using gradient descent: $\theta_{i}\leftarrow\theta_{i}-\eta_{\theta_{i}}\frac{\partial L_{\text{total}}}{\partial\theta_{i}}$ and $\mathbf{z}_{i}\leftarrow\mathbf{z}_{i}-\eta_{z}\frac{\partial L_{\text{total}}}{\partial\mathbf{z}_{i}}$. By optimizing $\mathbf{z}$, the modulation embedding learns to adjust its influence to minimize the loss, effectively enhancing inter-agent coordination and ensuring that each agent can dynamically adjust its output according to the state of the preceding agent.
The complete training process, including the initialization of the modulation embedding, the forward pass through the agents, and the optimization of both model parameters and the modulation embedding, is detailed in Algorithm 1. This algorithm encapsulates our self-modulated fine-tuning approach, providing a comprehensive framework for improving the performance and coordination of our multi-agent video generation system.
Input: Initial agents’ parameters $\{\theta_{i}\}$; initial modulation embeddings $\{\mathbf{z}_{i}=\mathcal{E}_{i}(\text{"{[Mod]}"})\}_{i=1}^{n}$; number of iterations $N$; number of samples per iteration $S$; prompt set $P$; batch size $B$; number of epochs per iteration $K$; learning rates $\eta_{\theta_{i}}$, $\eta_{z_{i}}$
Output: Trained model parameters $\{\theta_{i}\}$
1 for iteration $n=1$ to $N$ do
2
3      Construct training dataset $\mathcal{D}_{n}$ using the data-free training strategy (see Sec. 3.2 );
4
5      for epoch $e=1$ to $K$ do
6             foreach batch $\{(P^{(b)},V^{(b)}_{\text{target}})\}_{b=1}^{B}$ in $\mathcal{D}_{n}$ do
7                   Compute text embeddings $\mathbf{e}_{i}^{(b)}=\mathcal{E}_{i}(P^{(b)})$;
8                   for $i=1$ to $n$ do
9                         Concatenate agent-specific modulation embedding with text embedding for all examples: $\tilde{\mathbf{e}}_{i}^{(b)}=[\mathbf{e}_{i}^{(b)};\mathbf{z}_{i}]$;
10                         Generate outputs for the batch: $O_{i}^{(b)}=\mathcal{M}_{i}(O_{i-1}^{(b)},\tilde{\mathbf{e}}_{i}^{(b)})$, where $O_{0}^{(b)}=\emptyset$;
11
12                  Compute the final loss for the batch: $L_{\text{final}}=\frac{1}{B}\sum_{b=1}^{B}L(O_{n}^{(b)},V^{(b)}_{\text{target}})$;
13
14                  for $i=1$ to $n$ do
15                         Update agent-specific modulation embedding: $\mathbf{z}_{i}\leftarrow\mathbf{z}_{i}-\eta_{z_{i}}\frac{\partial L_{\text{final}}}{\partial\mathbf{z}_{i}}$;
16                         Compute modulation factor: $||\mathbf{z}_{i}||_{2}=\sqrt{\sum_{k}z_{i,k}^{2}}$;
17                         Compute modulation factor for model $i$: $\alpha_{i}=||\mathbf{z}_{i}||_{2}/n$;
18                         Update model parameters with dynamic learning rate: $\theta_{i}\leftarrow\theta_{i}-\alpha_{i}\eta_{\theta_{i}}\frac{\partial L_{\text{final}}}{\partial\theta_{i}}$;
19
20
21
22
23return $\{\theta_{i}\}$;
Algorithm 1 Self-Modulated Multi-Agent Video Generation with Data-free Fine-tuning
Multimodal LLM Selection with Human-in-the-loop Control.
Despite the availability of numerous open-source video datasets, their quality varies significantly, making it challenging for the pretrained agents we use to effectively leverage these datasets to improve video generation performance. Also, manually filtering high-quality data is time-consuming and inefficient. To address this issue, we introduce a multimodal LLM data selection procedure with human-in-the-loop control. We first sample a batch of candidates from the agent system. Based on the generated candidates, we further leverage strong multimodal LLMs that support multi-frame and multi-video inputs to provide evaluation for the candidate set. When multiple multimodal LLMs agree on the best video from a set, we directly include it in the training dataset, the evaluation prompt examples detailed in Appendix A.6. However, when the LLMs’ evaluations differ, we introduce human reviewers for secondary screening. In such cases, human reviewers evaluate the generated videos and select the one with the highest quality for inclusion in the training set. If the reviewers determine that none of the candidates meet the quality standard, the entire set is discarded. This approach enhances the stability and robustness of the filtering process by combining LLMs’ automated assessments with human judgment, ensuring the model can handle complex or ambiguous cases.
Data-free Training Strategy.
During training, instead of using open-source datasets directly, we adopt a data-free strategy to synthesize training data dynamically. This method addresses challenges like inconsistent quality in open-source datasets and the lack of intermediate outputs required for training different agents. Given a user input prompt set $P=\{P^{(1)},P^{(2)},\dots,P^{(S)}\}$, we first utilize the LLM to generate an enhanced set of diverse prompts $P_{\text{enh}}=\{P_{\text{enh}}^{(1)},P_{\text{enh}}^{(2)},\dots,P_{\text{enh}}^{(S)}\}$. For each enhanced prompt $P_{\text{enh}}^{(s)}$, our initial workflow—comprising non-finetuned models with parameters $\{\theta_{i}\}_{i=1}^{M}$—is applied to generate a set of candidate videos $\mathcal{C}_{s}=\{V_{1},V_{2},V_{3},V_{4}\}$. Next, we use a human-in-the-loop process, in combination with multimodal LLM selection, to select the highest quality video $\hat{V}^{(s)}$ from the candidate set $\mathcal{C}_{s}$ for each prompt $P_{\text{enh}}^{(s)}$. The selected video $\hat{V}^{(s)}$ and corresponding prompt $P_{\text{enh}}^{(s)}$ form the training dataset for iteration $n$, denoted by $\mathcal{D}_{n}=\{(P_{\text{enh}}^{(s)},\hat{V}^{(s)})\}_{s=1}^{S}$. Using the constructed dataset $\mathcal{D}_{n}$, self-modulated multi-agent fine-tuning is performed to update the agent parameters $\{\theta_{i}\}_{i=1}^{M}$, where $M$ is the number of agents involved. The newly fine-tuned model parameters $\{\theta_{i}\}_{i=1}^{M}$ then become the starting point for the next round of data generation and fine-tuning, enabling iterative improvement over $N$ iterations. Thus, after $N$ iterations, the final optimized agent parameters $\{\theta_{i}\}$ are obtained.
Figure 2: Samples for text-to-video generation of Mora. Our approach can generate high-resolution, temporally consistent videos from text prompts. The samples shown are 480p resolution over 12 seconds duration at 276 frames in total.
Figure 3: Performance variations of Task 5 and Task 6 across different self-training iterations.
@@ 4.1 Setup
Tasks. We evaluated our proposed Mora framework on six diverse video generation tasks to comprehensively assess its capabilities. The six tasks are as follows: (1) Text-to-Video Generation, where videos are generated directly from textual prompts; (2) Image-to-Video Generation, which involves creating videos conditioned on both an input image and a textual description; (3) Extend Generated Videos, aiming to extend existing videos to produce longer sequences while maintaining content consistency; (4) Video-to-Video Editing, which edits existing videos based on textual instructions to produce modified versions; (5) Connect Videos, focusing on seamlessly connecting two videos to
Figure 4: Performance variations of Task 1 to Task 4 across different self-training iterations.
create a longer continuous sequence; and (6) Simulate Digital Worlds, where videos that simulate digital or virtual environments are generated from textual prompts.
Data.   For the text-to-video generation task, we utilized textual prompts inspired by those provided in the official Sora technical report (OpenAI, 2024b). To expand our dataset, we employed GPT-4 (OpenAI, 2023) in both few-shot and zero-shot settings to generate additional prompts. These prompts were then used as inputs for the text-to-video models to generate videos. For the other tasks, we used relevant input data appropriate for each task, such as images, videos, or textual instructions. For comparison with Sora, we utilized videos featured on its official website and technical report.
@@
Baseline.  In the text-to-video generation process, we selected and compared a wide range of open-source and commercial models, including Sora (OpenAI, 2024a), VideoCrafter1 (Chen et al., 2023a), MedelScope (Mod, ), Show-1 (Zhang et al., 2023a), Pika (pik, ), LaVie and Lavie-Interpolation (Wang et al., 2023a), Gen-2 (Gen, a), and CogVideo (Hong et al., 2022a). In the other five tasks, we compare Mora with Sora. More visual comparisons and detailed analysis of the 5 tasks can be found in the Appendix A.3.
Metrics.  We employed a combination of standard and self-defined metrics to evaluate performance across the six tasks. For text-to-video generation, we employed several metrics from Vbench (Huang et al., 2024) for comprehensive evaluation from two aspects: video quality and video condition consistency. For video quality measurement, we use six metrics: ❶ Object Consistency, ❷ Background Consistency, ❸ Motion Smoothness, ❹ Aesthetic Score, ❺ Dynamic Degree and ❻ Imaging Quality. For measuring video condition consistency, we use two metrics: ❶ Temporal Style and ❷ Appearance Style. When evaluating other tasks, due to the lack of quantitative metrics, we design the following four metrics: ❶ Video-Text Integration $VideoTI$, ❷ Temporal Consistency $TCON$, ❸ Temporal coherence $Tmean$ and ❹ Video Length. For the details of these metrics, please refer to the Appendix A.2.
4.2 Training
In the training process, We initially used GPT-4o to generate 16 different text prompts and directly produced four videos for each prompt using an untrained model. During the data selection phase, we utilized two Multimodal LLMs, LLaVA-OneVision (Li et al., 2024a) and Qwen2-VL (Wang et al., 2024), to evaluate each set of videos and select the best-performing sample for each prompt. In the training loop, newly generated text prompts from GPT-4o were used, and after the selection process, the best videos were fed into the trained model for further optimization, and a total of 96 data were generated. Figures 4 and Figure 3 illustrate the impact of training on the performance of different tasks. It is evident that, during the early iterations, all tasks exhibit significant performance improvements, though the rate of improvement gradually slows as more iterations are completed. For details on the hyper-parameters settings in these training, please refer to Appendix A.9.